# Distorting Embedding Space for Safety: A Defense Mechanism for Adversarially Robust Diffusion Models

### v1

@@ Appendix G Analysis of Target Safe Vector Generation Phase
Highest Similarity 48.75 17.02 26.32
Random Safe Vector 40.64 16.61 25.82
Lowest Similarity 34.07 17.25 25.75
Target Vector w/ $s_{g}=50$ 18.63 16.16 25.97
Target Vector w/ $s_{g}=100$ 8.32 15.72 26.00
Target Vector w/ $s_{g}=150$ 1.54 15.73 25.53
Target Vector w/ $s_{g}=200$ 0.64 15.44 25.52
Target Vector w/ $s_{g}=250$ 0.59 15.91 25.22
Target Vector w/ $s_{g}=300$ 0.36 16.82 24.87
Target Vector w/ $s_{g}=350$ 0.49 16.77 24.86
Appendix H Effect of Loss Coefficient $\lambda$
The coefficient $\lambda$ balances UEN and SEP to achieve effective defense against unsafe image generation while maintaining benign image generation quality. We explore the optimal $\lambda$ by varying its value, as shown in Table H.1. When $\lambda=0.0,0.1$, DES focuses on distorting the unsafe embedding space, achieving low ASRs (0.18% and 0.44%), but significantly compromises benign image quality with high FID (113.38 and 58.53), and low CLIP score (10.00 and 17.41). Higher values ($\lambda=0.4,0.5,0.6$) improve FID and CLIP scores but increase ASR. While $\lambda=0.2$ achieves the best ASR with a slight impact on FID and CLIP scores, $\lambda=0.3$ provides excellent FID and CLIP score while maintaining a low ASR of 0.64%. We select $\lambda=0.3$ as our default setting, though $\lambda=0.2$ can be an alternative when prioritizing defense performance, and $\lambda=0.4,0.5$ are suitable for focusing on benign image quality.
Table H.1: Performance analysis with varying coefficient $\lambda$.
@@ Appendix H Effect of Loss Coefficient $\lambda$
DES ($\lambda=0.0$) 0.18 113.38 10.00
DES ($\lambda=0.1$) 0.44 58.53 17.41
DES ($\lambda=0.2$) 0.35 18.77 24.21
DES ($\lambda=0.3$) 0.64 15.44 25.52
DES ($\lambda=0.4$) 2.04 14.96 26.11
DES ($\lambda=0.5$) 2.50 14.97 26.13
DES ($\lambda=0.6$) 8.54 15.23 26.41
Appendix I Effective Target of $\lambda$
In this paper, we treat $\lambda$ as a ratio of SEP while controlling the combined effect of UEN and NEN ($\mathcal{L}_{u}+\mathcal{L}_{n}$) with $1-\lambda$. While this approach effectively balances defense capability and benign image generation quality, we explore alternative configurations for controlling UEN, SEP, and NEN. As shown in Table I.1, we evaluate three different loss combinations: SEP+UEN ($\mathcal{L}_{s}+\mathcal{L}_{u}$), SEP+NEN ($\mathcal{L}_{s}+\mathcal{L}_{n}$), and SEP ($\mathcal{L}_{s}$), with $\lambda$ ranging from 0.1 to 0.5.
The SEP+UEN configuration ($\lambda(\mathcal{L}_{s}+\mathcal{L}_{u})+(1-\lambda)\mathcal{L}_{n}$) achieves high-quality benign image generation but exhibits higher ASRs (2.77-6.83%), indicating that combining SEP with UEN compromises defense capability. The SEP+NEN configuration ($\lambda(\mathcal{L}_{s}+\mathcal{L}_{n})+(1-\lambda)\mathcal{L}_{u}$) achieves the lowest ASRs but struggles with generation quality at lower $\lambda$ values, though it shows promising results at $\lambda=0.3$ and $0.4$. The SEP configuration ($\lambda\mathcal{L}_{s}+(1-\lambda)(\mathcal{L}_{u}+\mathcal{L}_{n})$) demonstrates the best balance between defense capability and generation quality, particularly at $\lambda=0.3$ where it achieves a low ASR (0.64%) while maintaining competitive FID (15.44) and CLIP scores (25.52). Based on these results, we adopt $\lambda=0.3$ with the SEP configuration in our implementation.
Table I.1: Performance analysis with varying target of $\lambda$.
@@ Appendix I Effective Target of $\lambda$
$\lambda(\mathcal{L}_{s}+\mathcal{L}_{u})+(1-\lambda)\mathcal{L}_{n}$ $\lambda=0.1$ 6.83 15.18 26.34
$\lambda=0.2$ 2.77 14.95 26.30
$\lambda=0.3$ 3.46 14.90 26.28
$\lambda=0.4$ 4.17 15.09 26.27
$\lambda=0.5$ 6.49 15.13 26.32
$\lambda(\mathcal{L}_{s}+\mathcal{L}_{n})+(1-\lambda)\mathcal{L}_{u}$ $\lambda=0.1$ 0.14 43.96 18.78
$\lambda=0.2$ 0.42 18.30 23.98
$\lambda=0.3$ 1.06 15.15 25.66
$\lambda=0.4$ 1.28 15.24 25.64
$\lambda=0.5$ 4.36 15.03 26.21
$\lambda\mathcal{L}_{s}+(1-\lambda)(\mathcal{L}_{u}+\mathcal{L}_{n})$ $\lambda=0.1$ 0.44 58.53 17.41
$\lambda=0.2$ 0.35 18.77 24.21
$\lambda=0.3$ 0.64 15.44 25.52
$\lambda=0.4$ 2.04 14.96 26.11
$\lambda=0.5$ 2.50 14.97 26.13
Appendix J Analysis of Embedding Space Distortion
@@
We visualize the embedding space distortion of adversarial prompts from SneakyPrompt, I2P, Ring-A-Bell, and P4D in Figure J.1. Our analysis demonstrates that DES successfully transforms the majority of adversarial embeddings into the safe embedding region while preserving the original positions of safe embeddings. We observe that some adversarial embeddings, particularly from I2P and SneakyPrompt, maintain their original positions. This phenomenon can be attributed to the distinct characteristics of these attack methods. The I2P dataset contains a mixture of safe and unsafe prompts, with some prompts showing 0.0% nudity percentage (Schramowski et al., 2023), explaining the observed mixed distribution and selective transformation of embeddings. SneakyPrompt, on the other hand, specifically constructs unsafe prompts that closely resemble safe prompts to bypass filtering-based defenses (Yang et al., 2024c). However, as evidenced by the relatively low ASRs for both SDv1.4 and SDv1.5 in Table 1, these prompts may not consistently generate unsafe content, which explains their partial transformation in the embedding space.
@@
Figure J.1: Embedding space visualization. t-SNE visualization demonstrates how DES transforms adversarial prompt embeddings toward safe regions while preserving safe embedding positions.
Appendix K Limitations
@@
While DES demonstrates strong performance in defending against NSFW attacks, we acknowledge several limitations that warrant future research. First, as DES focuses on text encoder modification, it primarily addresses text-based attacks. Image-based attacks would require complementary defense methods specifically designed for image components. Second, while our approach effectively defends open-source models like Stable Diffusion, closed-source models may not directly benefit from DES. Although the insights from our study could inform the development of their defense mechanisms, the plug-and-play nature of our DES-trained text encoder may not be directly applicable to closed-source systems.

### v2

@@ C.3 Analysis of Target Safe Vector Generation Phase
Highest Similarity 54.13 17.02 26.32
Random Safe Vector 44.39 16.61 25.82
Lowest Similarity 38.50 17.25 25.75
Target Vector w/ $\alpha=50$ 20.74 16.16 25.97
Target Vector w/ $\alpha=100$ 9.10 15.72 26.00
Target Vector w/ $\alpha=150$ 1.48 15.73 25.53
Target Vector w/ $\alpha=200$ 0.52 15.44 25.52
Target Vector w/ $\alpha=250$ 0.43 15.91 25.22
Target Vector w/ $\alpha=300$ 0.30 16.82 24.87
Target Vector w/ $\alpha=350$ 0.50 16.77 24.86
C.3.1 Effective Scaling Factor for Other Concepts
In the paper, the scaling factor $\alpha$ was first determined for the nudity concept, as detailed in Table 15, and this value was subsequently applied to experiments involving other concepts. In this section, we have conducted additional experiments on $\alpha$ for other NSFW concepts and Van Gogh concept. Our experiments suggest that the most effective $\alpha$ remains within a relatively close range, as shown in Table 16. For example, for other NSFW concepts, the best performance was observed when $\alpha\in 200,250$. For the Van Gogh concept, the range yielding the best results was slightly broader at $\alpha\in 150,200,250$. This indicates that the value of $\alpha$ does not vary significantly according to the target concept.
Table 16: Impact of scaling factors ($\alpha$) on other NSFW and Van Gogh concepts.
@@ C.3.1 Effective Scaling Factor for Other Concepts
0 31.35 11.14 12.55 26.47 15.29 33.18 21.66 25.63 16.96
50 16.40 5.36 6.93 9.86 8.01 15.07 10.27 25.57 17.15
100 9.79 3.30 3.03 4.24 4.37 10.40 5.86 25.49 17.64
150 6.22 1.65 2.16 2.00 4.00 6.54 3.76 25.14 18.35
200 4.23 1.10 0.87 0.50 1.33 3.27 1.88 24.90 19.10
250 3.31 1.10 0.43 0.87 1.46 3.27 1.74 24.41 19.86
300 1.98 0.28 0.00 0.50 1.21 1.87 0.97 22.73 25.55
[…]
0 12.0 25.99 17.15
50 6.0 26.08 16.91
100 6.0 26.11 16.70
150 2.0 26.06 16.75
200 2.0 26.08 16.67
250 2.0 26.05 16.64
300 4.0 26.03 16.64
C.4 Analysis of Embedding Space Distortion
@@
We visualize the embedding space distortion of adversarial prompts from SneakyPrompt, I2P, Ring-A-Bell, and P4D in Figure 19. Our analysis demonstrates that DES successfully transforms the majority of adversarial embeddings into the safe embedding region while preserving the original positions of safe embeddings. We observe that some adversarial embeddings, particularly from I2P and SneakyPrompt, maintain their original positions. This phenomenon can be attributed to the distinct characteristics of these attack methods. The I2P dataset contains a mixture of safe and unsafe prompts, with some prompts showing 0.0% nudity percentage [sld], explaining the observed mixed distribution and selective transformation of embeddings. SneakyPrompt, on the other hand, specifically constructs unsafe prompts that closely resemble safe prompts to bypass filtering-based defenses [sneaky]. However, as evidenced by the relatively low ASRs for both SDv1.5 and FLUX.1 in Table 2, these prompts may not consistently generate unsafe content, which explains their partial transformation in the embedding space.
@@
Figure 19: Embedding space visualization. t-SNE visualization demonstrates how DES transforms adversarial prompt embeddings toward safe regions while preserving safe embedding positions.
Appendix D Ablation Studies
D.1 Contributions of Each Loss Function
We analyze the contribution of each loss function through ablation studies, as shown in Table 17. Using only $\mathcal{L}_{u}$ achieves the lowest ASR (0.35%) but significantly degrades benign image generation quality (FID 106.34, CLIP score 9.63). $\mathcal{L}_{s}$ aligns safe vectors with their originals, substantially improving image quality (FID 15.77, CLIP score 25.07) with a slight ASR increase (1.01%). Incorporating $\mathcal{L}_{n}$ refines ASR to 0.52% by neutralizing the “nudity” embedding and further enhances image quality (FID 15.44, CLIP score 25.52), demonstrating the synergistic effect of the three loss components.
[…]
0.35 106.34 9.63
1.01 15.77 25.07
0.52 15.44 25.52
D.2 Effect of the Loss Adjustment Technique
Table 18 demonstrates the effect of the loss adjustment technique by comparing ASR, FID, and CLIP score when $\mathcal{L}_{s}$ is implemented with and without the loss adjustment. The absence of the loss adjustment results in the deterioration of all metrics, highlighting its role in enhancing the ability of $\mathcal{L}_{s}$ to preserve the safe embedding region while effectively handling safe embeddings in ambiguous regions.
Table 18: Analysis of the loss adjustment technique. Results demonstrate the contribution of the loss adjustment technique within $\mathcal{L}_{s}$.
[…]
w/o Adjustment 1.76 15.65 25.43
w/ Adjustment 0.52 15.44 25.52
D.3 Effect of Loss Coefficient $\lambda$
The coefficient $\lambda$ balances unsafe and safe losses to achieve effective defense against unsafe image generation while maintaining benign image generation quality. We explore the optimal $\lambda$ by varying its value, as shown in Table 19. When $\lambda=0.0,0.1$, DES focuses on distorting the unsafe embedding space, achieving low ASRs (0.14% and 0.38%), but significantly compromises benign image quality with high FID (113.38 and 58.53), and low CLIP score (10.00 and 17.41). Higher values ($\lambda=0.4,0.5,0.6$) improve FID and CLIP scores but increase ASR. While $\lambda=0.2$ achieves the best ASR with a slight impact on FID and CLIP scores, $\lambda=0.3$ provides excellent FID and CLIP score while maintaining a low ASR of 0.52%. We select $\lambda=0.3$ as our default setting, though $\lambda=0.2$ can be an alternative when prioritizing defense performance, and $\lambda=0.4,0.5$ are suitable for focusing on benign image quality.
Table 19: Performance analysis with varying coefficient $\lambda$.
[…]
DES ($\lambda=0.0$) 0.14 113.38 10.00
DES ($\lambda=0.1$) 0.38 58.53 17.41
DES ($\lambda=0.2$) 0.18 18.77 24.21
DES ($\lambda=0.3$) 0.52 15.44 25.52
DES ($\lambda=0.4$) 1.98 14.96 26.11
DES ($\lambda=0.5$) 2.55 14.97 26.13
DES ($\lambda=0.6$) 9.70 15.23 26.41
D.4 Effective Target of $\lambda$
In this paper, we treat $\lambda$ as a ratio of the safe loss while controlling the combined effect of unsafe and nudity neutralization losses ($\mathcal{L}_{u}+\mathcal{L}_{n}$) with $1-\lambda$. While this approach effectively balances defense capability and benign image generation quality, we explore alternative configurations for controlling unsafe, safe, and nudity neutralization losses. As shown in Table 20, we evaluate three different loss combinations: safe loss + unsafe loss ($\mathcal{L}_{s}+\mathcal{L}_{u}$), safe loss + nudity neutralization loss ($\mathcal{L}_{s}+\mathcal{L}_{n}$), and safe loss ($\mathcal{L}_{s}$), with $\lambda$ ranging from 0.1 to 0.5.
The safe loss + unsafe loss configuration ($\lambda(\mathcal{L}_{s}+\mathcal{L}_{u})+(1-\lambda)\mathcal{L}_{n}$) achieves high-quality benign image generation but exhibits higher ASRs (2.70-7.30%), indicating that combining safe loss with unsafe loss compromises defense capability. The safe loss + nudity neutralization loss configuration ($\lambda(\mathcal{L}_{s}+\mathcal{L}_{n})+(1-\lambda)\mathcal{L}_{u}$) achieves the lowest ASRs but struggles with generation quality at lower $\lambda$ values, though it shows promising results at $\lambda=0.3$ and $0.4$. The safe loss configuration ($\lambda\mathcal{L}_{s}+(1-\lambda)(\mathcal{L}_{u}+\mathcal{L}_{n})$) demonstrates the best balance between defense capability and generation quality, particularly at $\lambda=0.3$ where it achieves a low ASR (0.52%) while maintaining competitive FID (15.44) and CLIP scores (25.52). Based on these results, we adopt $\lambda=0.3$ with the safe loss configuration in our implementation.
Table 20: Performance analysis with varying target of $\lambda$.
[…]
$\lambda(\mathcal{L}_{s}+\mathcal{L}_{u})+(1-\lambda)\mathcal{L}_{n}$ $\lambda=0.1$ 7.54 15.18 26.34
$\lambda=0.2$ 2.70 14.95 26.30
$\lambda=0.3$ 3.56 14.90 26.28
$\lambda=0.4$ 4.44 15.09 26.27
$\lambda=0.5$ 7.30 15.13 26.32
$\lambda(\mathcal{L}_{s}+\mathcal{L}_{n})+(1-\lambda)\mathcal{L}_{u}$ $\lambda=0.1$ 0.13 43.96 18.78
$\lambda=0.2$ 0.26 18.30 23.98
$\lambda=0.3$ 0.95 15.15 25.66
$\lambda=0.4$ 1.25 15.24 25.64
$\lambda=0.5$ 4.84 15.03 26.21
$\lambda\mathcal{L}_{s}+(1-\lambda)(\mathcal{L}_{u}+\mathcal{L}_{n})$ $\lambda=0.1$ 0.38 58.53 17.41
$\lambda=0.2$ 0.18 18.77 24.21
$\lambda=0.3$ 0.52 15.44 25.52
$\lambda=0.4$ 1.98 14.96 26.11
$\lambda=0.5$ 2.55 14.97 26.13
Appendix E Remarks
E.1 CLIP Score and FID for Prompts Closer to “Nudity” Embedding
Since DES modifies the embedding space to suppress unsafe content, it may affect prompts that are semantically close to the “nudity” embedding. If so, the FID and CLIP Score for such prompts could degrade, whereas prompts that are far from these unsafe regions might remain unaffected. To analyze this, we computed the FID and CLIP Score on the COCO dataset for the top 25% of prompts most similar to the “nudity” concept, as well as the bottom 25% (i.e., most dissimilar). As shown in the tables below, both groups experience a slight reduction in CLIP Score, but FID improves (i.e., lower), suggesting better visual quality but slightly reduced text-image alignment.
Table 21: FID and CLIP Score for the top 25% of prompts most similar to the “nudity” embedding.
[…]
Before DES 36.85 26.59
After DES 35.12 25.75
Table 22: FID and CLIP Score for the bottom 25% of prompts most dissimilar to the “nudity” embedding.
[…]
Before DES 32.55 26.26
After DES 31.34 25.27
We also examined individual prompt cases. For example, the prompt “A shirtless man in a hat making lunch” has high semantic similarity to the nudity concept. The CLIP Score for this prompt dropped from 30.34 (before DES) to 22.06 (after DES). Similarly, for the prompt “The man is walking down the street with no shirt on,” the CLIP Score dropped from 28.49 to 20.45. However, even prompts far from the nudity concept also show slight decreases. For instance, “Street signs, corner of Lynn and Bigelow. Taken 11.01.2009 23:58.” showed a drop from 21.47 to 20.33, and “A school bus and a silver car waiting at a railroad crossing for a train to go past.” dropped from 28.91 to 24.34. These results suggest that while DES does introduce some semantic distortion even to safe prompts, it generally preserves overall visual quality and remains consistent across both close and distant semantic regions.
E.2 Cultural Bias
We acknowledge cultural differences in defining sexual content. DES includes tunable parameters ($\lambda,\alpha$) to adjust suppression strength, allowing sensitivity calibration of the system. For example, decreasing $\lambda$ or $\alpha$ may allow for milder content while still preventing explicit sexual content. This flexibility allows DES to be adapted to different cultural or regulatory standards.
E.3 Limitations
@@
While DES demonstrates strong performance in mitigating sexual content generation, we acknowledge several limitations that warrant future research. First, as DES focuses on text encoder modification, it primarily addresses text-based attacks. Image-based attacks would require complementary defense methods specifically designed for image components. Second, while our approach effectively defends open-source models like Stable Diffusion, closed-source models may not directly benefit from DES. Although the insights from our study could inform the development of their defense mechanisms, our DES-trained text encoder may not be directly applicable to closed-source systems.
Appendix F Broader Impacts
Our work addresses the challenge of defending T2I diffusion models against sexual content generation. As these models become widely available, preventing misuse while maintaining functionality is important. DES provides a practical defense solution that effectively prevents sexual content generation while preserving the model’s ability to generate high-quality images. The positive impacts include improved safety in AI-generated content and reduced potential for model misuse.